# Rigidity aspects of a cosmological singularity theorem

### v2

@@ II. $f$ double covers a one-sided surface $K$
. We can go to the double cover $\tilde{V}$ such that the lift $\tilde{K}$ of $K$ is a connected, orientable, separating, incompressible embedded minimal surface. Thus by our previous arguments $\tilde{V}$ has a semibundle structure. Furthermore by construction both components of $\tilde{V}\setminus\tilde{K}$ are diffeomorphic to $V\setminus K$. But then $V$ is diffeomorphic to $(\tilde{K}\tilde{\times}I\setminus\tilde{K})\cup K$, i.e. it itself has a semibundle structure (in particular $\tilde{K}$ is totally geodesic in $V$).
@@
We first show that $V$ has to be prime. Let $\tilde{V}$ be its orientable double cover. Apply Theorem 1 to the corresponding covering spacetime $\tilde{M}$. Therefore $\tilde{V}$ is finitely covered by a surface bundle over $\mathbb{S}^{1}$ – call it $\tilde{V}^{\prime}$. We see that $\tilde{V}^{\prime}$ is prime: If the genus of the fibers is $\geq 1$, then its universal cover is $\mathbb{R}^{3}$ and hence irreducible. This contradicts the assumed reducibility of $V$, [21, Prop. 1.6]. If the genus is zero, then $\tilde{V}^{\prime}$ is homeomorphic to $\mathbb{S}^{1}\times\mathbb{S}^{2}$ and hence prime.
To finish the argument, we use the following fact [12, p. 7], which we prove below. Fact: The only nonprime manifold covered by a prime manifold is $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$ double covered by $\mathbb{S}^{1}\times\mathbb{S}^{2}$.
@@
Then $V$ contains an embedded 2-sided projective plane $\mathbb{RP}^{2}$ by the projective plane theorem (see [22, Thm. 1.1] or [20, Thm. 4.12]). It follows that there exists an embedded stable least area surface $\Sigma$ homeomorphic to $\mathbb{RP}^{2}$ [23, Prop. 2.3]. Moreover, $\Sigma$ is necessarily 2-sided [24, Lem. 2]. Note that $0\neq[\Sigma]\in\pi_{2}(V)$ since $\Sigma$ is incompressible in $V$ by Proposition 2.1 in [23]. By Lemma 2.1 in [25] the fundamental group of the orientable double cover $\tilde{V}$ of $V$ is torsion-free. Since $\tilde{V}$ is reducible (by the sphere theorem), part I of this proof shows that $\tilde{V}$ is either $\mathbb{S}^{1}\times\mathbb{S}^{2}$ or $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$ but only the former has torsion-free fundamental group. Furthermore [25, Lem. 2.1] also implies that $V$ is homotopic to $V^{\prime}:=\mathbb{S}^{1}\times\mathbb{RP}^{2}$.
Since $\pi_{2}(V^{\prime})\cong\pi_{2}(\mathbb{RP}^{2})$ there is a homotopy from $\Sigma$ in $V$ to a nonseparating embedded fiber $\Sigma^{\prime}\cong\mathbb{RP}^{2}$ in $V^{\prime}$. Therefore there is a nontrivial loop in $c^{\prime}$ in $V^{\prime}$ which intersects $\Sigma^{\prime}$ transversally only once. We remark that transversality is a generic condition and thus by the homotopy from $V$ to $V^{\prime}$, this loop is homotopic to a nontrivial loop $c$ in $V$ which intersects $\Sigma$ transversally only once (mod 2) (see also Chapters 2.3 and 2.4 in [26], in particular exercise 2 in section 2.4). Thus $\Sigma$ is nonseparating.
@@
We now prove the claim. First, since $V$ is nonorientable, by [20, Lemmas 6.6 and 6.7] it contains a two-sided, incompressible, nonseparating embedding $g\colon\Sigma\to V$. Moreover, $V$ contains no two-sided $\mathbb{RP}^{2}$ since $\pi_{2}(V)=0$; therefore it’s $P^{2}$-irreducible and hence a Haken manifold. There exists a two-sided incompressible least area immersion $f\colon\Sigma\to V$ homotopic to $g$ [15, Thm. 3.1]. Then either $f$ is an embedding or it double covers a one-sided surface $K$ and $g(\Sigma)$ bounds a submanifold of $V$ which is a twisted $I$-bundle over a surface isotopic to $K$ [15, Thm. 5.1]. However, the latter case contradicts the fact that $g$ is nonseparating. Lastly, since $g$ is nonseparating and $f$ is homotopic to $g$, it follows that $f$ is also nonseparating since the intersection number mod 2 is preserved under homotopy (see the corollary on page 79 of [26]).
@@
This implies that there is an embedding $g\colon\mathbb{S}^{2}\to V$ which separates and generates a nontrivial element of $\pi_{2}(V)$ [21, Prop. 3.7 and 3.10]. Then, by [27, Thm. 7] and [28, Thm. 5.8], $g$ is homotopic to a non-trivial minimal immersion $f\colon\mathbb{S}^{2}\to V$ which is either embedded or double covers an embedded projective plane. In the former case we can apply the same arguments as in the previous results to obtain a totally geodesic fibration within $V$ where the fibers are MOTS/MITS within $M$. In the latter case we can lift $f$ to a double cover $V^{\prime}$, which is also homeomorphic to $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$, to find a minimal embedding $f^{\prime}\colon\mathbb{S}^{2}\to V$ homotopic to $f$.
@@ 4 Discussion, Examples
$ds^{2}$ $=$ $-dt^{2}+\frac{3}{\Lambda}\cosh^{2}(\sqrt{\frac{\Lambda}{3}}t)[d\tau^{2}+\sin^{2}\tau(d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2})]$ (4.1)
@@
${\cal K}=\frac{1}{2}\frac{d}{dt}g=g\sqrt{\frac{\Lambda}{3}}\cosh\sqrt{\frac{\Lambda}{3}}t\leavevmode\nobreak\ \sinh\sqrt{\frac{\Lambda}{3}}t.$ (4.2)
@@
The $t=\mathrm{const.}$ surfaces are diffeomorphic to $\mathbb{S}^{3}$, which is parametrized by the Euler angles $\psi\in[0,4\pi);\leavevmode\nobreak\ \theta\in[0,\pi),\leavevmode\nobreak\ \phi\in[0,2\pi)$. The "Taub-region" $U>0$ is given by $m-\sqrt{m^{2}+\ell^{2}}<t<m+\sqrt{m^{2}+\ell^{2}}.$
@@
$ds^{2}$ $=$ $-dt^{2}+\frac{1}{\Lambda}\left(\cosh^{2}\sqrt{\Lambda}t\,d\alpha^{2}+d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2}\right)$
$t\in(-\infty,\infty)\quad\alpha\in[0,L),\leavevmode\nobreak\ L\in\mathbb{R},\quad\theta\in[0,\pi)\quad\phi\in[0,2\pi).$
@@
${\cal K}_{\alpha\alpha}=\frac{1}{2}\frac{d}{dt}g_{\alpha\alpha}={\Lambda}^{-1/2}\cosh\sqrt{\Lambda}t\leavevmode\nobreak\ \sinh\sqrt{\Lambda}t.$ (4.4)
@@
Fajman and Kröncke [32] have analyzed spacetimes where the spatial sections are products of Einstein spaces. In $3$ spatial dimensions their ansatz takes the form [32, Thm. 1.6]
$ds^{2}=-dt^{2}+\frac{b(t)^{2}}{\Lambda}d\alpha^{2}+\frac{a(t)^{2}}{\widehat{\Lambda}}[d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2}]$ (4.5)
@@
${\cal K}=\frac{1}{2}\frac{d}{dt}g={\widehat{\Lambda}}^{-1}\,\frac{da}{dt}\,\mbox{diag}[0,1,\sin^{2}\vartheta]$ (4.6)
is strictly $2$-convex. (The degeneracy at $\vartheta=0$ is of course an artifact of the polar coordinates.) Since this $t=0$ section of (4.5) is obviously not a spherical space, Theorem 0 yields null geodesic incompleteness which is consistent with the findings of Thm. 1.6. in [32]. In this example Theorems 1 and 2 are inconclusive as (4.5) is also consistent with consequences (iii) and (iv).
@@
$ds^{2}=r^{2}\left(\frac{dr^{2}}{\sigma}+d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2}\right)=r(\alpha,m,\Lambda)^{2}\left(d\alpha^{2}+d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2}\right).$ (4.7)
@@
We have to require that the period $P(m,\Lambda)$ defined below "fits" on the Nariai torus, which restricts its size $L$ via
$P(m,\Lambda)=2\int_{r_{b}}^{r_{c}}\sigma^{-\frac{1}{2}}dr=\frac{L}{j}\qquad(j\in\mathbb{N}).$ (4.9)
Thus the Cauchy surface contains $j$ pairs of black hole and cosmological horizons, and (4.9) entails a relation between the parameters $m$, $L$ and $j$. As is well-known and suggested by the "horizon" terminology, the spacetime is (future and past) null geodesically incomplete. Thus we are in case (i) of Theorems 1 and 2; the other cases are ruled out since $\partial/\partial\alpha$ is not Killing and the foliation is not totally geodesic.
@@
As in Example 5 we consider data for Einstein’s equations with $\Lambda>0$ conformal to the Nariai torus but now with conformal factor depending on $\vartheta$ rather than $\alpha$, viz.
$ds^{2}=\varphi(\vartheta,J,\Lambda)^{4}\left(d\alpha^{2}+d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2}\right).$ (4.10)
@@
Each of these solutions $\varphi_{J}$ now leads to a solution of the constraints $(V,g,{\cal K})$ which (due to their stability) share the $U(1)\times U(1)$ symmetry of the coefficients in the Lichnerowicz equation (4.13), cf. e.g. [34, Prop. 2]. This implies that the metric indeed takes the form (4.10), which is a totally geodesic foliation. Moreover, it is obvious from (4.14) that the data are both ($t$-$\phi$)-symmetric as well as $(t-\alpha)$ symmetric as defined above. The constant $J$ is the "Komar" angular momentum of the data with respect to the Killing field $\partial/\partial\phi$.
@@
$ds^{2}=\rho^{2}\left(\frac{dr^{2}}{\sigma}+\frac{d\vartheta^{2}}{\chi}\right)+\frac{\sin^{2}\vartheta}{\kappa^{2}\rho^{2}}\left[\chi(r^{2}+a^{2})^{2}-\sigma a^{2}\sin^{2}\vartheta\right]d\phi^{2},$ (4.15)
@@
$\sigma=(r^{2}+a^{2})\left(1-\frac{\Lambda r^{2}}{3}\right)-2mr,\quad\rho^{2}=r^{2}+a^{2}\cos^{2}\vartheta,\quad\chi=1+\frac{\Lambda a^{2}\cos^{2}\vartheta}{3}\,.$ (4.16)
@@
$P(m,a,\Lambda)=2\int_{r_{b}}^{r_{c}}\sigma^{-\frac{1}{2}}dr=\frac{L}{j}\qquad j\in\mathbb{N}$ (4.17)
@@
Let $V$ be a closed flat three-dimensional manifold equipped with time-symmetric vacuum initial data for the Einstein equations. The corresponding maximal globally hyperbolic vacuum development is the Lorentzian product $\mathbb{R}\times V$, which is clearly geodesically complete. There are 10 such data sets of the form $V=\mathbb{E}^{3}/\Gamma$ where $\Gamma$ is a discrete and fixed-point free symmetry group of the Euclidean space $\mathbb{E}^{3}$. There are six orientable ones $G_{1},\dotsc,G_{6}$ and four non-orientable ones $B_{1},\dotsc,B_{4}$ (see the classification in [43].) For each one Theorem 0 does not apply, but Theorem 1 does. If $V$ is either the flat torus $G_{1}=\mathbb{T}^{3}$ or one of other 4 possibilities with $H_{2}(V,\mathbb{Z})\neq 0$, then Theorem 1 (iii) applies directly to $V$ (without going to a cover). On the other hand, $G_{6}$ (i.e., the Hantzsche–Wendt manifold), has $H_{2}(G_{6},\mathbb{Z})=0$ and is not a surface bundle over the circle. Now Theorem 1 (iii) still applies to a cover. Indeed, if one applies the construction in the proof of Theorem 1 to $G_{6}$, then one obtains its double cover $G_{2}$. However, as mentioned in the introduction, Proposition 1 applies directly to $G_{6}$ (i.e. without invoking a cover).
@@
Let $(V,g,{\cal K})$ be an initial data set for the vacuum Einstein equations with or without a cosmological constant. Assume $g$ is a hyperbolic metric on a closed 3-manifold $V$ and $\cal K$ is $2$-convex. Then Theorem 1 implies that the resulting maximal globally hyperbolic development is past null geodesically incomplete. For if this was not the case, then there would be a three-manifold with a hyperbolic metric which is a surface bundle over $\mathbb{S}^{1}$ containing totally geodesic fibers. But this is impossible; see [44, Thm. 3.3] and the discussion on mapping tori below it.
@@
Acknowledgments
We are grateful to Greg Galloway for fruitful interaction in all stages of the present work, most importantly in connection with the virtual positive $b_{1}$-conjecture (cf. Remark 5). We also acknowledge helpful correspondence with Florian Beyer, Shuli Chen, Klaus Kröncke, Vanderson Lima and Franco Vargas Pallete.
Eric Ling was supported by Carlsberg Foundation CF21-0680 and Danmarks Grundforskningsfond CPH-GEOTOP-DNRF151. Carl Rossdeutscher and Walter Simon were funded by the Austrian Science Fund (FWF) [Grant DOI 10.55776/P35078]; Roland Steinbauer was funded in part by FWF as well [Grant DOI 10.55776/EFP6]. Part of the research on this paper was performed while two of the authors were participants of the program “New Frontiers in Curvature: Flows, General Relativity, Minimal Submanifolds, and Symmetry” at the Simons Laufer Mathematical Sciences Institute (SLMath). We thank SLMath for its support. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v3

@@ II. $f$ double covers a one-sided surface $K$
. We can go to the double cover $\tilde{V}$ such that the lift $\tilde{K}$ of $K$ is a connected, orientable, separating, incompressible embedded minimal surface. Thus, by our previous arguments used in I, $\tilde{V}$ has a semibundle structure. Furthermore by construction both components of $\tilde{V}\setminus\tilde{K}$ are diffeomorphic to $V\setminus K$. But then $V$ is diffeomorphic to $(\tilde{K}\tilde{\times}I\setminus\tilde{K})\cup K$, i.e. it itself has a semibundle structure (in particular $\tilde{K}$ is totally geodesic in $V$).
@@
We first show that $V$ has to be prime. Let $\tilde{V}$ be its orientable double cover. Apply Theorem 1 to the corresponding covering spacetime $\tilde{M}$. Therefore $\tilde{V}$ is finitely covered by a surface bundle over $\mathbb{S}^{1}$ – call it $\tilde{V}^{\prime}$. We see that $\tilde{V}^{\prime}$ is prime: If the genus of the fibers is $\geq 1$, then its universal cover is $\mathbb{R}^{3}$ and hence irreducible. This contradicts the assumed reducibility of $V$, [22, Prop. 1.6]. If the genus is zero, then $\tilde{V}^{\prime}$ is homeomorphic to $\mathbb{S}^{1}\times\mathbb{S}^{2}$ and hence prime.
To finish the argument, we use the following fact [23, p. 7], which we prove below. Fact: The only non-prime manifold covered by a prime manifold is $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$ double covered by $\mathbb{S}^{1}\times\mathbb{S}^{2}$.
@@
Then $V$ contains an embedded 2-sided projective plane $\mathbb{RP}^{2}$ by the projective plane theorem (see [13, Thm. 1.1] or [26, Thm. 4.12]). It follows that there exists an embedded stable least area surface $\Sigma$ homeomorphic to $\mathbb{RP}^{2}$ [11, Prop. 2.3]. Moreover, $\Sigma$ is necessarily 2-sided [25, Lem. 2]. Note that $0\neq[\Sigma]\in\pi_{2}(V)$ since $\Sigma$ is incompressible in $V$ by Proposition 2.1 in [11]. By Lemma 2.1 in [41] the fundamental group of the orientable double cover $\tilde{V}$ of $V$ is torsion-free. Since $\tilde{V}$ is reducible (by the sphere theorem), part I of this proof shows that $\tilde{V}$ is either $\mathbb{S}^{1}\times\mathbb{S}^{2}$ or $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$ but only the former has torsion-free fundamental group. Furthermore [41, Lem. 2.1] also implies that $V$ is homotopic to $V^{\prime}:=\mathbb{S}^{1}\times\mathbb{RP}^{2}$.
Since $\pi_{2}(V^{\prime})\cong\pi_{2}(\mathbb{RP}^{2})$ there is a homotopy from $\Sigma$ in $V$ to a nonseparating embedded fiber $\Sigma^{\prime}\cong\mathbb{RP}^{2}$ in $V^{\prime}$. Therefore there is a non-trivial loop in $c^{\prime}$ in $V^{\prime}$ which intersects $\Sigma^{\prime}$ transversally only once. We remark that transversality is a generic condition and thus by the homotopy from $V$ to $V^{\prime}$, this loop is homotopic to a non-trivial loop $c$ in $V$ which intersects $\Sigma$ transversally only once (mod 2) (see also Chapters 2.3 and 2.4 in [19], in particular exercise 2 in section 2.4). Thus $\Sigma$ is nonseparating.
@@
We now prove the claim. First, since $V$ is nonorientable, by [26, Lemmas 6.6 and 6.7] it contains a two-sided, incompressible, nonseparating embedding $g\colon\Sigma\to V$. Moreover, $V$ contains no two-sided $\mathbb{RP}^{2}$ since $\pi_{2}(V)=0$; therefore it’s $P^{2}$-irreducible and hence a Haken manifold. There exists a two-sided incompressible least area immersion $f\colon\Sigma\to V$ homotopic to $g$ [21, Thm. 3.1]. Then either $f$ is an embedding or it double covers a one-sided surface $K$ and $g(\Sigma)$ bounds a submanifold of $V$ which is a twisted $I$-bundle over a surface isotopic to $K$ [21, Thm. 5.1]. However, the latter case contradicts the fact that $g$ is nonseparating. Lastly, since $g$ is nonseparating and $f$ is homotopic to $g$, it follows that $f$ is also nonseparating since the intersection number mod 2 is preserved under homotopy (see the corollary on page 79 of [19]).
@@
This implies that there is an embedding $g\colon\mathbb{S}^{2}\to V$ which separates and generates a non-trivial element of $\pi_{2}(V)$ [22, Prop. 3.7 and 3.10]. Then, by [30, Thm. 7] and [38, Thm. 5.8], $g$ is homotopic to a non-trivial minimal immersion $f\colon\mathbb{S}^{2}\to V$ which is either embedded or double covers an embedded projective plane. In the former case we can apply the same arguments as in the previous results to obtain a totally geodesic fibration within $V$ where the fibers are MOTS/MITS within $M$. In the latter case we can lift $f$ to a double cover $V^{\prime}$, which is also homeomorphic to $\mathbb{RP}^{3}\#\mathbb{RP}^{3}$, to find a minimal embedding $f^{\prime}\colon\mathbb{S}^{2}\to V$ homotopic to $f$.
@@ 4 Discussion, Examples
$ds^{2}$ $=$ $-dt^{2}+\frac{3}{\Lambda}\cosh^{2}\left(\sqrt{\frac{\Lambda}{3}}t\right)[d\tau^{2}+\sin^{2}\tau(d\theta^{2}+\sin^{2}\theta d\phi^{2})]$ (4.1)
@@
${\cal K}=\frac{1}{2}\frac{d}{dt}g=\sqrt{\frac{\Lambda}{3}}\tanh\left(\sqrt{\frac{\Lambda}{3}}t\right)g.$ (4.2)
@@
The $t=\mathrm{const.}$ surfaces are diffeomorphic to $\mathbb{S}^{3}$, which is parametrized by the Euler angles $\psi\in[0,4\pi);\penalty 10000\ \theta\in[0,\pi),\penalty 10000\ \phi\in[0,2\pi)$. The "Taub-region" $U>0$ is given by $m-\sqrt{m^{2}+\ell^{2}}<t<m+\sqrt{m^{2}+\ell^{2}}.$
@@
$ds^{2}$ $=$ $-dt^{2}+\frac{1}{\Lambda}\left(\cosh^{2}\sqrt{\Lambda}t\,d\alpha^{2}+d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$
$t\in(-\infty,\infty)\quad\alpha\in[0,L),\penalty 10000\ L\in\mathbb{R},\quad\theta\in[0,\pi)\quad\phi\in[0,2\pi).$
@@
${\cal K}_{\alpha\alpha}=\frac{1}{2}\frac{d}{dt}g_{\alpha\alpha}={\Lambda}^{-1/2}\cosh\sqrt{\Lambda}t\penalty 10000\ \sinh\sqrt{\Lambda}t.$ (4.4)
@@
Fajman and Kröncke [14] have analyzed spacetimes where the spatial sections are products of Einstein spaces. In $3$ spatial dimensions their ansatz takes the form [14, Thm. 1.6]
$ds^{2}=-dt^{2}+\frac{b(t)^{2}}{\Lambda}d\alpha^{2}+\frac{a(t)^{2}}{\widehat{\Lambda}}[d\theta^{2}+\sin^{2}\theta d\phi^{2}]$ (4.5)
@@
${\cal K}=\frac{1}{2}\frac{d}{dt}g={\widehat{\Lambda}}^{-1}\,\frac{da}{dt}\,\mbox{diag}[0,1,\sin^{2}\theta]$ (4.6)
is strictly $2$-convex. (The degeneracy at $\theta=0$ is of course an artifact of the polar coordinates.) Since this $t=0$ section of (4.5) is obviously not a spherical space, Theorem 0 yields null geodesic incompleteness which is consistent with the findings of Thm. 1.6. in [14]. In this example Theorems 1 and 2 are inconclusive as (4.5) is also consistent with consequences (iii) and (iv).
@@
$ds^{2}=r^{2}\left(\frac{dr^{2}}{\sigma}+d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)=r(\alpha,m,\Lambda)^{2}\left(d\alpha^{2}+d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (4.7)
@@
The following equation in which we introduce the length $L$ of the Nariai torus and the period $P(m,\Lambda)$ states that the period has to "fit" on the torus, i.e.
$L=\int_{0}^{L}d\alpha=2j\int_{r_{b}}^{r_{c}}\sigma^{-\frac{1}{2}}dr=jP(m,\Lambda)$ (4.9)
for some $j\in\mathbb{N}$. Thus the Cauchy surface contains $j$ pairs of black hole and cosmological horizons, and (4.9) entails a relation between the parameters $m$, $L$ and $j$. As is well-known and suggested by the "horizon" terminology, the spacetime is (future and past) null geodesically incomplete. Thus we are in case (i) of Theorems 1 and 2; the other cases are ruled out since $\partial/\partial\alpha$ is not Killing and the foliation is not totally geodesic.
@@
As in Example 5 we consider data for Einstein’s equations with $\Lambda>0$ conformal to the Nariai torus but now with conformal factor depending on $\theta$ rather than $\alpha$, viz.
$ds^{2}=\varphi(\theta,J,\Lambda)^{4}\left(d\alpha^{2}+d\vartheta^{2}+\sin^{2}\vartheta d\phi^{2}\right).$ (4.10)
@@
Each of these solutions $\varphi_{J}$ now leads to a solution of the constraints $(V,g,{\cal K})$ which (due to their stability) share the $U(1)\times U(1)$ symmetry of the coefficients in the Lichnerowicz equation (4.13), cf. e.g. [10, Prop. 2]. This implies that the metric indeed takes the form (4.10), which is a totally geodesic foliation. Moreover, it is obvious from (4.14) that the data are both ($t$-$\phi$)-symmetric as well as $(t-\alpha)$ symmetric as defined above. The constant $J$ is the "Komar" angular momentum of the data with respect to the Killing field $\partial/\partial\phi$.
@@
$ds^{2}=\rho^{2}\left(\frac{dr^{2}}{\sigma}+\frac{d\theta^{2}}{\chi}\right)+\frac{\sin^{2}\theta}{\kappa^{2}\rho^{2}}\left[\chi(r^{2}+a^{2})^{2}-\sigma a^{2}\sin^{2}\theta\right]d\phi^{2},$ (4.15)
@@
$\sigma=(r^{2}+a^{2})\left(1-\frac{\Lambda r^{2}}{3}\right)-2mr,\quad\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\quad\chi=1+\frac{\Lambda a^{2}\cos^{2}\theta}{3}\,.$ (4.16)
@@
$L=\int_{0}^{L}d\alpha=2j\int_{r_{b}}^{r_{c}}\sigma^{-\frac{1}{2}}dr=jP(m,a,\Lambda)\qquad j\in\mathbb{N}.$ (4.17)
@@
Let $V$ be a closed flat three-dimensional manifold equipped with time-symmetric vacuum initial data for the Einstein equations. The corresponding maximal globally hyperbolic vacuum development is the Lorentzian product $\mathbb{R}\times V$, which is clearly geodesically complete. There are 10 such data sets of the form $V=\mathbb{E}^{3}/\Gamma$ where $\Gamma$ is a discrete and fixed-point free symmetry group of the Euclidean space $\mathbb{E}^{3}$. There are six orientable ones $G_{1},\dotsc,G_{6}$ and four nonorientable ones $B_{1},\dotsc,B_{4}$ (see the classification in [45].) For each one Theorem 0 does not apply, but Theorem 1 does. If $V$ is either the flat torus $G_{1}=\mathbb{T}^{3}$ or one of other 4 possibilities with $H_{2}(V,\mathbb{Z})\neq 0$, then Theorem 1 (iii) applies directly to $V$ (without going to a cover). On the other hand, $G_{6}$ (i.e., the Hantzsche–Wendt manifold), has $H_{2}(G_{6},\mathbb{Z})=0$ and is not a surface bundle over the circle. Now Theorem 1 (iii) still applies to a cover. Indeed, if one applies the construction in the proof of Theorem 1 to $G_{6}$, then one obtains its double cover $G_{2}$. However, as mentioned in the introduction, Proposition 1 applies directly to $G_{6}$ (i.e. without invoking a cover).
@@
Let $(V,g,{\cal K})$ be an initial data set for the vacuum Einstein equations with or without a cosmological constant. Assume $g$ is a hyperbolic metric on a closed 3-manifold $V$ and $\cal K$ is $2$-convex. Then Theorem 1 implies that the resulting maximal globally hyperbolic development is past null geodesically incomplete. For if this was not the case, then there would be a three-manifold with a hyperbolic metric which is a surface bundle over $\mathbb{S}^{1}$ containing totally geodesic fibers. But this is impossible; see [43, Thm. 3.3] and the discussion on mapping tori below it.
@@
Acknowledgments.
We are grateful to Greg Galloway for fruitful interaction in all stages of the present work, most importantly in connection with the virtual positive $b_{1}$-conjecture (cf. Remark 5). We also acknowledge helpful correspondence with Florian Beyer, Shuli Chen, Klaus Kröncke, Vanderson Lima and Franco Vargas Pallete. Furthermore, we thank the referee for useful comments.
Eric Ling was supported by Carlsberg Foundation CF21-0680 and Danmarks Grundforskningsfond CPH-GEOTOP-DNRF151. Carl Rossdeutscher and Walter Simon were funded by the Austrian Science Fund (FWF) [Grant DOI 10.55776/ P35078]; Roland Steinbauer was funded in part by FWF as well [Grant DOI 10.55776/EFP6]. Part of the research on this paper was performed while two of the authors were participants of the program “New Frontiers in Curvature: Flows, General Relativity, Minimal Submanifolds, and Symmetry” at the Simons Laufer Mathematical Sciences Institute (SLMath). We thank SLMath for its support. We also acknowledge the support of the Simons Center for Geometry and Physics, where parts of this work were carried out during the program “Geometry and Convergence in Mathematical General Relativity”. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.